OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2)

OpenStudy (anonymous):

y = - x - 6 y = - x - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

@jessica_the_meow

OpenStudy (anonymous):